# Python-Based Quantum Chemistry Calculations with GPU Acceleration

### v1

@@ 4.1 Compatibility with PySCF
Listing 1: Example of PySCF with GPU support
@@ 4.2 Accelerated Functionalities
Dispersion corrections are usually much cheaper than DFT self-consistent iterations. And these contributions are calculated with dftd3/simple-dftd3 and dftd4/dftd4 packages without GPU acceleration. Although DFT-D3 has been implemented in Torch and Jax, these implementations will introduce the dependencies of heavy-weight machine learning framework. For robustness, we build simple-dftd3 and dftd4 in GPU4PySCF, and disable OpenMP. Nonlocal corrections are carefully optimized at CUDA level due to its expensive computational cost.
PCM/SMD solvent models require solving Poisson-Boltzmann equation in each SCF iteration. The computational complexity of solving Poisson-Boltzmann equation is cubic-scaling with respect to the number of grids on molecular surface. Although it is cheaper than the quartic-scaling hybrid DFTs, it would be the bottleneck without GPU acceleration. To address this, we have re-engineered the PCM/SMD solvent models for GPU execution using Python, recognizing the potential need for further optimizations to minimize the memory requirements.
ECP contributions are calculated only once in SCF. We reuse PySCF routines on CPU to calculate the ECP contributions in SCF, gradient, and Hessian. Other contributions of DFT are still accelerated with GPU for a molecular system with ECP. The computational cost is acceptable if the system only has a few heavy atoms. We leave this as future work to further accelerate ECP calculations.
@@
4.3 Coordinating with other packages
The development of PySCF/GPU4PySCF benefits from the contributions from the open-source community. GPU4PySCF, acting as an extension of PySCF, can be seamlessly incorporated with the open-source packages compatible with PySCF. We highlight the roles of the following fundamental packages working with PySCF/GPU4PySCF.
  • BSE [14]: basis set exchange for newly developed basis set.
  • GeomeTRIC [77]: geometric optimization and transition state search.
  • LibXC [15]: DFT exchange correlation functionals.
  • DFTD3 [99]/DFTD4 [100]: dispersion corrections.
  • Libcint [101]: fundamental GTO integrals.
  • CuPy [102]: A drop-in replacement of Numpy & SciPy for GPU.
@@ 5 Limitations and Future development
Quantum chemistry is a field characterized by its complex algorithms and significant computational demands. Accelerated computing technologies, such as GPUs, offer promising ways to tackle these challenges, yet their adoption in quantum chemistry has been slower than in other areas. Although the package delivers significant speedup for various tasks, we admit our implementations are not the optimal. We recognize the current limitations and suggest focused areas for future development to enhance computational efficiency and capabilities.
@@
Direct SCF & Other Integral Schemes. The Direct Self-Consistent Field (SCF) method has been particularly explored for GPU acceleration due to its minimal memory requirements and scalability with system size, especially when screening techniques are applied. The efficiency of Direct SCF is largely dependent on the integral computation strategy. We utilize Rys quadrature due to its simplicity and versatility in calculating Gaussian Type Orbital (GTO) integrals on GPUs. In the current implementations, this integral scheme is applied up to g functions. While it may not be the fastest, it simplifies the integral program for various operators. The selection of the optimal algorithm for integral computations is crucial for maximizing performance.
@@ Acknowledgements
We thank Dr. John Herbert from The Ohio State University for providing the implementation details of CHELPG charge in Q-Chem 6.1.
@@ Appendix C Accuracy for Different Basis Sets
Table 9: 6-31G.

### v2

@@ 4.1 Improved Compatibility with PySCF
Listing 1: Example of PySCF with GPU support. The code changes from CPU to GPU are colored in red.
With the new design, GPU4PySCF v1.0 has been meticulously engineered to maintain a high degree of compatibility with PySCF, ensuring that any settings related to accuracy in PySCF are seamlessly applicable in GPU4PySCF. For example, consider DFT calculations.
@@ 4.2 Accelerated Functionalities
Dispersion corrections are usually much cheaper than DFT self-consistent iterations. And these contributions are calculated with dftd3/simple-dftd3 and dftd4/dftd4 packages without GPU acceleration. Although DFT-D3 has been implemented in Torch and Jax, these implementations will introduce the dependencies of heavy-weight machine learning framework. For robustness, we rebuild simple-dftd3 and dftd4, and disable OpenMP. The same interfaces are used in both PySCF and GPU4PySCF. Nonlocal corrections are carefully optimized at CUDA level due to its expensive computational cost.
PCM/SMD solvent models require solving Poisson-Boltzmann equation in each SCF iteration. The computational complexity of solving Poisson-Boltzmann equation is cubic-scaling with respect to the number of grids on molecular surface. Although it is cheaper than the quartic-scaling hybrid DFTs, it would be the bottleneck without GPU acceleration. To address this, we have re-engineered the PCM/SMD solvent models for GPU execution using Python, recognizing the potential need for further optimizations to minimize the memory requirements. To maintain the compatibility with PySCF we also introduced PCM/SMD solvent models in PySCF v2.4 and added nuclear Hessian in PySCF v2.6.
ECP contributions are calculated only once in SCF. We reuse PySCF routines on CPU to calculate the ECP contributions in SCF, gradient, and Hessian. Other contributions of DFT are still accelerated with GPU for a molecular system with ECP. The computational cost is acceptable if the system only has a few heavy atoms. We leave the GPU acceleration of ECP as future work.
@@
4.3 Improved Interfaces to External Libraries
One of the advantages that we have benefited from by working with PySCF and GPU4PySCF is that we can benefit from the contributions from the open-source community. In particular, GPU4PySCF, acting as an extension of PySCF, can be seamlessly incorporated with the open-source packages compatible with PySCF. In addition to the many libraries that PySCF depends on BSE [14] (basis set exchange for newly developed basis set), GeomeTRIC [79] (geometric optimization and transition state search), DFTD3 [100]/DFTD4 [101] (dispersion corrections) and Libcint [102] (fundamental GTO integrals), our work in GPU4PySCF has benefited extensively from the CUDA ecosystem:
  • CuPy [103], a drop-in replacement for NumPy & SciPy for GPU.
  • CUDA LibXC. CUDA support is still experimental in the latest version of libXC (v6.2). In this work, the exchange-correlation potentials and their derivatives are evaluated on the GPU using a customized Python API.
  • cuTENSOR. cuTENSOR is used as the default tensor contraction engine. It takes the advantage of tensor cores, and provides convenient tensor operations on NVIDIA GPUs. Other tensor contraction packages, such as opt_einsum [104] and cuquantum [105] are also supported.
@@ 5 Limitations and Future Development
Quantum chemistry is a field characterized by its complex algorithms and significant computational demands. Accelerated computing technologies, such as GPUs, offer promising ways to tackle these challenges, yet their adoption in quantum chemistry has been slower than in other areas. With our many contributions to the GPU4PySCF package, we can now deliver significant speedup for various tasks. As demonstrated through our applications, GPU4PySCF can now be applied in many different practical settings, bringing the full power of GPUs to bear on a large part of quantum chemistry modeling. We can thus recommend GPU4PySCF as a powerful module for the PySCF ecosystem in industrial applications.
At the same time, our implementations are certainly not optimal and there are many current limitations. As dedicated stakeholders in this project, we outline our vision on some of the important future developments to further enhance the capabilities of this package.
@@
Direct SCF & Other Integral Schemes. The Direct Self-Consistent Field (SCF) method has been particularly explored for GPU acceleration due to its minimal memory requirements and scalability with system size, especially when screening techniques are applied. The efficiency of Direct SCF is largely dependent on the integral computation strategy. As described in Ref. [20] the start of the GPU4PySCF project was built around Direct SCF using a Rys quadrature scheme to compute the 4-center two-electron integrals. The Rys quadrature scheme was selected mainly because of its simplicity for implementing various operators and derivatives. In future work, it may be worth revisiting the integral scheme for direct SCF calculations.
@@
Low-level Integration with other PySCF-based Packages. Due to fundamental differences between the underlying CPU algorithms and their corresponding GPU algorithms, the low-level APIs of PySCF and GPU4PySCF are not fully compatible. As a result, packages relying on low-level PySCF APIs cannot be directly accelerated with GPU4PySCF. Additionally, we are aware of other groups working on PySCF-based GPU implementations. Other designs could use JAX or Torch as the linear algebra backend. Defining the interfaces between these packages remains an area of active exploration.
@@ Acknowledgements
The authors would like to thank the organizers of the 1st annual PySCF Developers Meeting. The initial results for GPU acceleration of PySCF were presented at that meeting by Dr. Qiming Sun before the public release of the GPU4PySCF repository. The authors thank all the attendees for their discussions. After the meeting, Dr. Garnet Chan and Dr. Qiming Sun initiated the public GPU4PySCF repository onGitHub, and the Caltech team developed the GPU accelerated 4-center GTO integrals and direct SCF scheme that formed the initial version of GPU4PySCF. This initial version is described further in Ref. [20]. The authors also would like to thank Dr. John Herbert from The Ohio State University for providing the implementation details of CHELPG charge in Q-Chem 6.1.
Author Contributions
Xiaojie Wu and Qiming Sun designed the framework and APIs in this work; built CI; analyzed the performance; conceptualized the paper and planned the application projects. Qiming Sun contributed this work when he worked as an independent researcher. Xiaojie Wu optimized the framework of DFT; implemented density fitting, gradient, Hessian, and solvent models; performed benchmarks of DF modules with calculating solvation free energy, transition state search. Zhichen Pu developed open-shell HF/DFT, CHELPG charge, and NMR. Tianze Zheng and Wen Yan performed the applications in torsion scan and dimer interaction energies. Wenzhi Ma performed the applications of Fukui functions and ESP. Zhichen Pu, Sheng Gong, Yumin Zhang and Weihao Gao studied Lithium ion solvation structure and electrochemical stability windows of electrolytes. Weiluo Ren and Xiang Li incorporated GPU4PySCF with neural networks. Yu Xia, Zhengxiao Wu, and Mian Huo contributed the engineering improvements. All authors contributed the paper draft.
@@ Appendix C Accuracy for Different Basis Sets
Table 9: Difference between GPU4PySCF and Q-Chem 6.1, using 6-31G basis.